OpenStudy (jt950):

4. The pressure of an ideal gas is doubled during a process in which the energy given up as heat by the gas equals the work done on the gas. As a result, the volume is: A) doubled B) halved C) unchanged D) need more information to answer E) nonsense, the process is impossible Please Explain.

OpenStudy (john_es):

By the first principle of thermodynamic, \[\Delta U=W+Q\] As Q (is negative because is given up) equals W, then the process is isotherm, so, \[W=Q\Rightarrow \Delta T=constant\]Then it is possible to apply, \[P_1V_1=P_2V_2\Rightarrow P_1V_1=2P_1V_2\Rightarrow V_2=V_1/2\] And the correct answer is b.

OpenStudy (jt950):

Thanks , When do we use Q=U+W and when do we use U=W+Q What difference between them? What exactly define Q , U , W?

OpenStudy (jt950):

@John_ES

OpenStudy (john_es):

It all dependes on the sign of the work and the heat transfered. The usual is to put always, \[\Delta U=Q+W\] And then, put the signs. For example, for a process that gives heat, Q-> -Q, and do work against the exterior W>0, so, \[\Delta U=-|Q|+W\Rightarrow \Delta U+|Q|=W\]